OpenStudy (anonymous):

The legs of a right triangle measure 3x and 15. If the hypotenuse measures 3x + 3, what is the value of x? a.221 b.12 c.36 d.16

OpenStudy (anonymous):

first get the formula for a right angled triangle then slot all the numbers in and then solve for x. first try that then we'll see how you do. I don't want to just give you the answer

OpenStudy (anonymous):

its called the pythagorus theorem: a^ + b^ = c^ a is one of the legs and so is b then c is the hypotenuse. when you are solving for the hypotenuse you"ll have something like (3x + 3)^2 then make sure to simplify it which would be 9x^2+18x+9

OpenStudy (anonymous):

pythagorus a^2+b^2=c^2

OpenStudy (anonymous):

I don't really understand...is the pythagorus theorem: a^ + b^ = c^ the formula for a right angled triangle?

OpenStudy (anonymous):

yes it is the same thing, its actually a squared + b squared = c squared or you can say x squared + y squared = r squared

OpenStudy (anonymous):

so here is the answer: a^2 + b^2 = c^2 since one leg is 15 and the other is 3x therefore its going to be like this (15)^2+(3x)^2 =(3x+3)^2 225+9x^2=9x^2+18x+9 take all the numbers to one side and all the x's to one side as well so you have 225-9=9x^2-9x^2+18x+9 which gives you: 216=18x when you solve for x you get x=12

OpenStudy (anonymous):

ohh i didnt understand that you were meaning squared in your formula thanks i get it now.. : )

OpenStudy (anonymous):

pleasure